### v1

# 61A-Bot: AI homework assistance in CS1 is fast and cheap—but is it helpful?

Abstract.
Chatbot interfaces for LLMs enable students to get immediate, interactive help on homework assignments, but even a thoughtfully-designed bot may not serve all pedagogical goals. In this paper, we report on the development and deployment of a GPT-4-based interactive homework assistant (“61A-Bot”) for students in a large CS1 course; over 2000 students made over 100,000 requests of our bot across two semesters. Our assistant offers one-shot, contextual feedback, through both a “Get Help” button within a popular code editor, as well as a “get feedback” feature within our command-line autograder. These triggers wrap student code in a custom prompt that supports our pedagogical goals and avoids providing solutions directly. We discuss our development process and deployment, then analyze possible impacts of our Bot on students, primarily through student feedback and how long it takes students to complete homework problems. We ask: how does access to 61A-Bot impact homework completion time and subsequent course performance? In addition to reductions in homework-related question rates in our course forum, we find substantial reductions in homework completion time. These are most pronounced for students in the percentile, with reductions of over 30 minutes, over 4 standard deviations faster than the mean in prior semesters. However, it is not clear that these effects transfer to assignment contexts where the Bot is not available: we observe speedups in some contexts, no change in others, and some assignments later in the semester even show a slowdown instead. Though we have begun to disentangle these effects, further research is needed.
1. Introduction
The recent wide availability of Large Language Models (LLMs) has given students in introductory Computer Science (CS) courses a tempting alternative to asking for help on programming assignments—and potentially waiting hours to receive it. However, while naively used LLMs do help students solve assigned problems, they typically do so by providing correct answers along with explanations, allowing students to avoid the process of developing solutions themselves and the learning associated with this process. A number of recent reports (Denny et al., 2023a, c; Finnie-Ansley et al., 2023; Hellas et al., 2023; Kazemitabaar et al., 2024; Liffiton et al., 2023; Wang et al., 2024) present more thoughtful approaches: new systems, also based on LLMs, geared towards offering guidance and assistance without providing direct solutions.
These systems succeed in providing guidance without direct solutions, and students report finding them helpful (Kazemitabaar et al., 2024; Liffiton et al., 2023)—but how are they helpful? What impact do these systems have on student performance when integrated into a traditional CS1 course? In this paper, we start to answer these questions: we first describe the assistant we developed and deployed to 2000 students in our CS1 course over the past year. Then, we report on a preliminary observational investigation of our assistant’s impact on those students, finding that our assistant does indeed help students complete homework more quickly, but does not necessarily help them learn the material better.
We begin by reporting on our experiences developing and deploying a “TA bot” homework assistant for students in the large CS1 course (CS 61A) we run at our large public university. We call our bot 61A-Bot. Students receive bot feedback in two ways: either (1) by clicking on a “Get Help” button in their code editor (see Figure 1), receiving back a pop-up notification containing advice for how to proceed on the homework assignment, given their current code; or (2) through a “get feedback” option to the existing course “autograder” command-line tool—allowing all students access to the same feedback tool, irrespective of the choice of code editor.
Our bot then identifies the particular assignment the student is working on, collects their code, and wraps these in a custom prompt for GPT-4. We designed this prompt to steer GPT-4 towards feedback that mirrors how we ourselves typically approach student questions, aligned with recent work in this area (Markel and Guo, 2021): identifying whether the student understands the question, which concepts students might need reinforcement on, and whether they have a plan, and then helping students by providing conceptual, debugging, or planning support as appropriate.
We then report on:
Initial findings of students’ subjective experiences based on student surveys over two semesters, and
A retrospective analysis of student homework completion patterns, looking for possible impacts our bot may have had on students’ homework experiences and overall performance.
Required by our Institutional Review Board (IRB) to avoid differential access, we did not engage in a randomized control experimental design. Instead, we rely on observational data, analyzing completion-time performance for homeworks (for which students had Bot access) and similar-in-scope labs and more complex projects (for which students did not have Bot access) in our one semester of complete deployment—comparing these with historical completion-time data from previous semesters. While this approach does not necessarily allow us to draw conclusive causal inferences, the associations we find are nonetheless suggestive and call for further inquiry.
In our analysis, we observe remarkably shorter homework completion times, after the deployment of our Bot, on comparable problems from previous semesters. Suggestive of a direct causal effect is our parallel observation that completion times for comparable labs, where the Bot is never available, show no similar shortening after our deployment. Meanwhile, completion times on projects, where the Bot is also not available, are somewhat shorter after Bot deployment, but exhibit a smaller reduction than the homeworks. Finally, we also observe weak evidence of what may be over-reliance or a dependency effect: performance on bot-never-available labs shows some performance degradation, and students report that they feel as though labs take much longer than they would if the bot were available. In line with these time reductions, we also find substantial reductions in the number of online forum questions about homework problems. We report on these observations in detail in the following sections, and conclude with implications and opportunities for future inquiry.
2. Background & Related Work
Generative models such as ChatGPT,111https://chat.openai.com/ OpenAI Codex (Chen et al., 2021), DeepMind AlphaCode (Li et al., 2022), Amazon CodeWhispere222https://aws.amazon.com/codewhisperer/, and GitHub Copilot333https://copilot.github.com/ offer promising opportunities for enriching the learning experience of students. These models have already been leveraged by educators in different areas of Computing education (Finnie-Ansley et al., 2022; Denny et al., 2023c; Hellas et al., 2023; Denny et al., 2023a), where they accelerate content generation and seem to be impacting the relevant skills students gain in introductory CS courses. Researchers have studied LLMs in areas such as generating code explanations (Leinonen et al., 2023a; Becker et al., 2023; MacNeil et al., 2023b; Donlevy, 2023), providing personalized immediate feedback, enhancing programming error messages (Leinonen et al., 2023b; Wang et al., 2024), generating discussion forum responses (Mitra et al., 2024; Liu et al., 2024), and automatic creation of personalized programming exercises and tutorials (Sarsa et al., 2022; Yuan et al., 2023; Reeves et al., 2023) to enhance the comprehensiveness of course materials.
However, the integration of LLMs in CS1 instruction comes with challenges. Students could become overly reliant on automation (a concern at least as old as calculators (Demana and Waits, 2000)), potentially hindering their development of critical problem-solving skills—though recent work suggests these negative effects can be avoided, at least for programming assistance (Kazemitabaar et al., 2023). Taken to an extreme, the resulting absence of human interaction could have negative effects, alongside other ethical concerns related to plagiarism and the responsible use of LLM-generated code. To maximize the benefits of LLMs while mitigating these challenges, a thoughtful and balanced approach to their incorporation into CS1 courses is essential (Denny et al., 2023b; Mirhosseini et al., 2023; MacNeil et al., 2023a).
Through deployments of LLMs as intelligent tutors, students can receive immediate, personalized support and guidance, which would ideally foster a deeper understanding of coding concepts and promote self-paced learning—just as with pre-LLM Intelligent Tutoring Systems (see (Crow et al., 2018) for a review, and (Suzuki et al., 2017) for an example). The ability of LLMs to generate tailored resources, such as new, personalized tutorials and newly-generated code examples, not only expands the available learning materials but also accommodates students’ varying learning styles and preferences—though these generated materials are not always better (Pardos and Bhandari, 2023). One recent RCT of an error message enhancement system (Wang et al., 2024) begins to shed some light on how different approaches to hint generation, including GPT-4-augmented error messages, lead to different patterns of student behavior—finding that GPT-4-augmented messages can lead to persistent reductions in the number of attempts it takes students to resolve a bug in their code compared to a typical compiler or runtime error message. Educators should integrate LLMs as complementary tools, striking a balance between automation and human interaction while emphasizing the development of critical problem-solving skills and responsible coding practices, ultimately serving students better in their CS education.
Researchers are also increasingly integrating LLM-based chatbots in courses (Donlevy, 2023; Wu et al., 2022) and online educational websites (Ofgang, 2023) to provide immediate personalized feedback, and in tools in supporting students’ development of programming skills (Prather et al., 2023; Finnie-Ansley et al., 2023; Cipriano and Alves, 2023). These include CodeHelp (Liffiton et al., 2023; Sheese et al., 2024) and CodeAid (Kazemitabaar et al., 2024), two systems (and deployments) that bear a number of similarities to our own—though those systems enable students to ask questions, while ours (we believe uniquely) builds on the student’s history of prior assistant hints and code changes in response. That body of work finds that students primarily seek answers to programming questions and debugging help; we focus on the latter here.
Our paper contributes to the body of educational research and pedagogical innovation, offering a preliminary evaluation of the impact of an LLM-based AI homework assistant on student performance on course assignments.
3. Design & Deployment
In designing our assistant, we chose to address one of the more challenging bottlenecks we faced in our large-scale course: the availability of tutors and other staff to help students when they get stuck on homework problems. In particular, we chose to focus primarily on the kinds of debugging assistance our staff are often asked for.
Three concerns—hallucinations, students sharing personal information with a third party, and the harms from an unmonitored chat interaction—led us to deploy a one-shot “Get Help” interaction mode without an opportunity for “chat” follow-up. This meant that one valuable pedagogical tool—having students explain their understanding of the problem—would remain out of reach in this initial deployment.
Following a common tutoring pattern (Markel and Guo, 2021), we designed a prompt that would try to assess student conceptual knowledge, based on the provided code, and offer syntactical, logical, or even template-code suggestions—but not solutions. This prompt explicitly includes a sequence of questions to consider in response to the student code:
Is the student missing conceptual knowledge?
Is their current code on the right track?
How close are they to a solution?
Were they able to follow previous advice?
Do they have a reasonable plan?
Though we avoided students explicitly writing natural language “chat” messages to the bot, we did want some degree of continuity—which we achieved by also including up to three prior (student code, bot advice) exchanges, if available (enabling question 4).
In addition to the steps above, the prompt also includes a per-problem instruction block, which we used for about 10% of problems, and more general instructions such as Do not give the student the answer or any code. and Limit your response to a sentence or two at most.
3.1. Development & Prompt Challenges
Based on the literature, we expected that engineering dialogues through an LLM would be challenging in ways that would be hard to address (Zamfirescu-Pereira et al., 2023), and indeed we found this to be the case. Early on, in testing extended dialogue interactions, we found an increasing likelihood that GPT-4 would provide a direct solution as the conversation extended—validating our decision to start with a single-shot request rather than dialogue. Many initial challenges were addressable through prompt changes alone: for example, by including the instruction “Don’t assume a problem needs to use recursion unless it’s explicit” we avoided spurious suggestions that loop-based code should be rewritten using recursion.
One particularly frustrating challenge illustrates a number of the failures we observed early on: an over-eagerness to “correct” student code that was, in fact, already correct. One such homework problem asked students to fill in template code to compute , using the operators add and sub based on whether (see Fig. 1). With an early version of our prompt, GPT-4 would consistently question the correct solution, asking students to think about what should be done in the case .
Hoping to find a prompt that would prove robust to variations in course content, we initially avoided any problem-specific text beyond a statement of the problem itself. However, we ultimately added a problem-specific note for those homework problems that required extra steering—in the example above, this note contains a solution to the problem.
3.2. Deployment
Students access our bot through two modalities. In one modality, students click on a “Get Help” button (see Figure 1) in the VS Code444Visual Studio Code, https://code.visualstudio.com editor toolbar to activate our 61A-Bot extension. The extension collects students’ code, makes a best guess of which homework problem the student is working on (as several problems often appear in a single file), and constructs a request. In the second modality, students run an autograder555 OK Client, https://github.com/okpy/ok-client in the command line which collects the student’s code, any errors from executing test cases, and constructs a request from these. These requests reach a server run by our instructional staff, which wraps the student data in our prompt, passes it to GPT-4, and logs the request and GPT-4’s response for further analysis.
Students are informed when installing the software that, in using our assistant, all code they write will be sent to OpenAI via Microsoft Azure (Microsoft, 2024), and that they should not include any content in their code files (e.g., comments) that they do not want to share.
3.2.1. Deployment Timeline
We piloted and continued development on 61A-Bot throughout the academic year 2023-2024. In Fall 2023, we deployed an initial pilot of 61A-Bot to an experimental section of 400 students. This was followed by a full-scale deployment in week 7 for the approximately 1300 students across two sections. In conjunction with the wider deployment, we also enabled access through the autograder tool students could already run from the command line to validate their code against a set of test cases. In Spring 2024, all 900 enrolled students had access to both modalities of the bot from the start of the semester. Our qualitative analysis thus draws on both semesters, while our quantitative analysis is focused on the full deployment in Spring 2024.
4. Findings
Students’ adoption of 61A-Bot was immediate, and usage exploded once we integrated access into our autograder. In this section, we detail student assessments of their experiences, followed by an investigation of the possible impacts of 61A-Bot on homework completion times, exam score distributions, and reliance on course staff. Unless otherwise noted, statistical accounts in this section come exclusively from Spring 2024, the semester in which our fully-developed Bot was deployed for all students (see §3.2.1 for deployment details).
4.1. Usage & Reliance on Staff
Usage patterns suggest that students are finding 61A-Bot helpful, returning to it multiple times as they engage in homework: across our pilot and full deployment semesters, over 2000 students made a total of 105,689 requests of our bot. The median student in our full deployment semester made 25 requests to our bot, rising to 80 requests for the student at the 95th percentile. As expected, usage increases as the assignment deadline nears and is concentrated in the late afternoons and evenings—a pattern similar to the usage reported in (Liffiton et al., 2023; Kazemitabaar et al., 2024)—with a peak request rate of 291 requests/hour.
Over 93% of 61A-Bot queries were made via the autograder instead of VS Code, a skew likely explained by the “get feedback?” query to students on every run of the command-line autograder, which historically students run liberally. By contrast, the “Get Help” button in VS Code requires a manual installation of an extension, in addition to the use of VS Code, which our course does not require.
This engagement correlates with a reduction in help requests on our online discussion forum for students to receive asynchronous help. There is a substantial (30%!) decrease in the number of questions asked per student throughout from Spring 2023 to Spring 2024, from 1.741 to 1.185 questions per student. The impact on homework-specific questions is even larger, showing a 75% decrease from 0.344 homework questions per student in SP23 to 0.088 in SP24—see Table 1. (Note the rough consistency in question counts per student in Fall 2022 and Fall 2023.)
| Fa22 | Sp23 | Fa23 | Sp24 | |
| # Students | 1656 | 1169 | 1407 | 872 |
| # Questions | 1853 | 2035 | 1823 | 1033 |
| per student | 1.119 | 1.741 | 1.296 | 1.185 |
| # HW Questions | 234 | 402 | 177 | 77 |
| per student | 0.141 | 0.344 | 0.126 | 0.088 |
4.2. Student Reception
We sought student feedback in two ways. First, we queried students for their assessment of each individual hint, which we received for approximately 27% () of queries.666These counts reflect queries from only those students who gave consent for their data to be used for research. Of these, 70% () were rated as “helpful,” with 45% of these, , reporting that the problem was now resolved. A further 10% () were rated as “not helpful, but made sense,” while the final 20% () were rated as insensible, misleading, or wrong.
Second, we formally surveyed students on their usage and perceptions of the Bot. In Fall 2023, we conducted a non-anonymous survey at the end of the semester to which 49% () of students responded. Students were asked to rate how much they used the bot and how helpful they found it on a scale from 1 to 5. As expected, those who reported more usage also found it more helpful.
In Spring 2024, we conducted a non-anonymous survey at the end of the semester, to which 89% () students responded. On a scale from 1 to 5, we asked students to rate their bot usage, bot helpfulness, bot reliability, and overall satisfaction with the bot. Finally, we asked them whether or not they recommend that the bot be available to students in future semesters. The results from these surveys can be found in Figure 2.
In the Spring 2024 survey, we also asked two optional questions: What did students like the most and least about learning with the bot? We include a representative response for each here: What students liked the most:
“What I loved about the bot is that it allowed me to get feedback when I didn’t have access to a tutor. Accordingly, instead of banging my head against the wall for hours, I was able to get feedback about what I was doing wrong and correct the mistakes. For me personally, I would have had a lot more success in this class if I would have had access to the bot for labs. (Labs on average took me about 2 and a half hours to complete, and sometimes longer if I didn’t have access to a tutor).”
What students liked the least:
“Sometimes the answers were slightly vague. Of course, the bot can’t simply spit out the answer, but sometimes it was frustrating how it would say ‘you’re on the right track, but there seems to be a conceptual misunderstanding with __’ — the explanations for the blank could be a bit jargon-filled and didn’t always directly help me resolve the misunderstanding due to imprecise language.”
Students generally appreciated 61A-Bot’s accessibility, debugging capabilities, and time savings. However, the hints were sometimes too vague for the students to make changes, while other times, the bot was too specific and gave away too much. Despite our attempts to steer GPT-4 towards rephrasing and a broader diversity of responses to repeat inputs, these incidents still occur.
4.3. Effects on Assignment Completion Time
Having established that students like 61A-Bot, and that it can provide hints that appear useful, we began seeking ways to quantitatively identify what impact—if any—the bot was having on student outcomes. This proved to be a challenging task, in part because the impact is nuanced and specific, as we detail in this section. As mentioned earlier, our IRB does not allow differential access to tools in courses, so we could not run a randomized control trial. Instead, here we compare student performance from our semester of full deployment, Spring 2024 (SP24), with performance from prior semesters, going back through Spring 2021 (SP21), comprising a total of 1,643,613 data points from 6,034 students.
Though the Bot is deployed for homework assignments it is not available for lab assignments (which are similar in scope, but intended to be completed in groups) nor for projects (which are multi-week and build on prior results). To the extent that performance differences on homework assignments between pre- and post-deployment semesters are inconsistent with performance differences on lab assignments and projects, some of this difference may be attributable to the use of the Bot. In our course, content is typically sequenced as follows: new material is first introduced in lectures, then practiced in labs, reinforced in individual homework problems, and then applied to practice in projects.
4.3.1. Data
Our course “autograder” records every attempt a student makes to test their code, storing a “snapshot” of student code on our instructional servers, along with a student identifier and a timestamp; we use these to reconstruct a coarse-grained history of student progress. Autograder use is unlimited, and students typically revise their programs repeatedly until they are ready to submit their final code. Because nearly all students submit code that successfully passes all test cases, the metric we consider here is time to completion, rather than passing test cases or other measures of code quality.
We calculate an approximate total completion time by summing the timestamp deltas between snapshots, recognizing that we may miss the time spent on the problem before the first snapshot. To account for students completing the homework across multiple sessions, we ignore time deltas above a 60-minute threshold—a value chosen somewhat arbitrarily, but we confirm that the results we report here are robust to values in the range of 20 to 120 minutes.777We conducted an independent analysis looking only at snapshot counts rather than summing time deltas. Those results were consistent with the findings we report here, as expected from (Piech et al., 2012); we report times here because we find the time metric to be more straightforward to reason about.
We compare Spring 2024 completion times only to prior Spring semesters, because Fall-semester students at our institution typically have greater preparation than Spring-semester students (due in part to how CS major requirements are structured). Because problem repetition across homeworks and labs has some variation across semesters, we select a subset of problems to maximizes consistency, and report these aggregations as, e.g., “HW 1”. Occasionally, one assignment in a particular semester differs sufficiently from other semesters that we omit it entirely.
4.3.2. Analytical Method
We examine the distribution of student assignment completion times using Cumulative Distribution Function (CDF) plots for specific homework assignments, labs, and projects. This CDF can be read as “What fraction of students (y-axis) complete the homework in less than some number of minutes (x-axis)?” Because not all semesters used exactly the same problems in each assignment, we analyze only that subset of problems that appear identically across most semesters in a given assignment, summing those problems’ completion times and reporting the sum as the total homework completion time for the assignment. (Some semesters are missing from some plots because of this variability in problem inclusion for assignments as mentioned earlier.) As with the choice of completion time vs. submission counts, we opted for this method of aggregation and reporting because we found it to be the clearest presentation of the effects we observed.888We performed the same analysis on individual problems on their own, as well as by comparing individual semesters with their prior (Spring) semesters and comparing across different subsets of problems and prior years; the results we report here are robust to all of the variations we tried, including many others that we do not report on here for space.
4.3.3. Results
Our primary finding is that student completion times on identical homework assignments are substantially faster in our post-deployment semester, Spring 2024, compared to prior semesters (3(a)). This effect is not seen in the similar lab assignments where the bot is not available (3(b)); a smaller, but still substantial, speedup is seen in projects (3(c)).
The data indicate that the Bot-available homework completion times in Spring 2024 are between faster among students in the middle percentile of completion times. This represents a reduction in time of Standard Deviations (SD) compared with completion times from prior Spring semesters, a large effect.
In contrast, no-Bot lab completion times fall within 1 SD of the prior semesters’ mean for the first few labs. Most later labs lack enough consistency across semesters for a robust comparison, but lab 8 does offer a glimpse into one possible outcome: here, students in Spring 2024 overall took several standard deviations more time than the mean for students in prior semesters.
Meanwhile, the no-Bot project completion times fall somewhere in between: Spring 2024 students completed these projects between faster than the pre-deployment mean—a speedup about half as large as the Bot-available homeworks.
5. Discussion
Overall, our findings are consistent with a causal link between the availability of our bot and faster homework completion times. There is no substantially different teaching methodology, set of course instructors, or set of course materials in Spring 2024 that could account for the specific reduction in homework completion times without a similar reduction in lab completion times—to our knowledge (as course instructors), 61A-Bot itself is the major difference in Spring 2024, and the major difference between homeworks on the one hand, and labs and projects on the other.
The speedup differences among homeworks, projects, and labs, could be due to a few causal factors. First, some of the benefits of using the Bot on homeworks could transfer to projects, but not labs—perhaps because projects around a given topic are assigned after students have already had the Bot’s help on the corresponding homework, but not on the corresponding labs, which are assigned before the homework.
Second, students in Spring 2024 may simply have stronger prior preparation in a way that affects projects, but not labs—though we are not sure what kind of preparation that could be. The consistency of performance on the first few labs in Spring 2024 and prior semesters suggests that Spring 2024 students do not differ substantially in their prior preparation for CS1.
The apparent gradual increase in lab completion times (starting in lab 4, continuing in lab 8, see 3(b)) are worth noting because they do not correlate with a corresponding increase in homework completion time, suggesting that the availability of the bot may be a cause. Students could be relying on the bot for certain activities that they might otherwise have learned more effectively to perform on their own, such as debugging code on previously unseen topics. We find some possible corroborating evidence for this effect in the survey results, where a top request for the course is to make the bot available for lab assignments too (see §4.2). But the improvement in project completion times suggests that if there is over-reliance on the Bot in the lab context, it does not extend to the project context—and it could be that lab 8 is merely an outlier in this regard, and not indicative of a robust pattern.
5.1. Limitations
We have been careful to identify limitations throughout this report but include this section here for completeness. Most importantly, our study is entirely observational: we had limited levers for randomization, and there could be uncontrolled causes for the effects we observe—for example, it could be that students are using ChatGPT for their homework anyway, despite the availability of 61A-Bot. We have tried to account for as many of these factors as possible through mechanism-oriented reasoning (e.g., why would they use ChatGPT only for homeworks, but not labs?), but we can not prove causality with our study design.
6. Conclusions
Our results suggest that 61A-Bot helps students complete homework more quickly, with oversized impacts for students who spent the most time on homework—a benefit that might even disproportionately support goals towards inclusion in CS. But, ideally, this type of scaffolding should recede over time as learners become more confident (Soloway et al., 1994). 61A-Bot has not yet achieved this goal.
However, the news is not entirely bad: whatever skills students may come to depend on the bot scaffolding for—whether debugging, problem-solving, or something else—are skills we can try to teach explicitly in our course, now that students spend less time on homework. Of course, it may turn out that the best way for students to learn those skills is to struggle more in debugging their homework assignments on their own, but we are skeptical of this view. Fortunately, effective homework assistants, like our 61A-Bot, can also help future researchers start to disentangle these effects, leading to a better understanding of the pedagogical value of struggling to debug code.
Guidelines around the inclusion of AI-based course materials and tools suggest that these should only be incorporated when we have a good understanding that their benefits outweigh their costs (Bala et al., 2023). We believe that 61A-Bot clears this bar, but that further research into improving outcomes and mitigating the costs we have started to expose is critical and urgent.
Acknowledgements.
This work was made possible by a few generous sources of support: an Inclusion Research Award from Google, and support for 61A-Bot’s use of Azure’s OpenAI API by Microsoft.References
- (1)
- Bala et al. (2023) Kavita Bala, Alex Colvin, Morten H. Christiansen, Allison Weiner Heinemann, Sarah Kreps, Lionel Levine, Christina Liang, David Mimno, Sasha Rush, Deirdre Snyder, Wendy E. Tarlow, Felix Thoemmes, Rob Vanderlan, Andrea Stevenson Won, Alan Zehnder, and Malte Ziewitz. 2023. Generative Artificial Intelligence for Education and Pedagogy | Center for Teaching Innovation. https://teaching.cornell.edu/generative-artificial-intelligence/cu-committee-report-generative-artificial-intelligence-education
- Becker et al. (2023) Brett A Becker, Paul Denny, James Finnie-Ansley, Andrew Luxton-Reilly, James Prather, and Eddie Antonio Santos. 2023. Programming Is Hard–Or at Least It Used to Be: Educational Opportunities And Challenges of AI Code Generation. (2023).
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. 2021. Evaluating large language models trained on code. arXiv preprint arXiv:2107.03374 (2021).
- Cipriano and Alves (2023) Bruno Pereira Cipriano and Pedro Alves. 2023. GPT-3 vs Object Oriented Programming Assignments: An Experience Report. In Proceedings of the 2023 Conference on Innovation and Technology in Computer Science Education V. 1. 61–67.
- Crow et al. (2018) Tyne Crow, Andrew Luxton-Reilly, and Burkhard Wuensche. 2018. Intelligent tutoring systems for programming education: a systematic review. In Proceedings of the 20th Australasian Computing Education Conference. 53–62.
- Demana and Waits (2000) Franklin Demana and BK Waits. 2000. Calculators in mathematics teaching and learning. Past, present, and future. In Learning Mathematics for a New Century (2000), 51–66.
- Denny et al. (2023a) Paul Denny, Brett A Becker, Juho Leinonen, and James Prather. 2023a. Chat Overflow: Artificially Intelligent Models for Computing Education-renAIssance or apocAIypse?. In Proceedings of the 2023 Conference on Innovation and Technology in Computer Science Education V. 1. 3–4.
- Denny et al. (2023b) Paul Denny, Juho Leinonen, James Prather, Andrew Luxton-Reilly, Thezyrie Amarouche, Brett A Becker, and Brent N Reeves. 2023b. Promptly: Using Prompt Problems to Teach Learners How to Effectively Utilize AI Code Generators. arXiv preprint arXiv:2307.16364 (2023).
- Denny et al. (2023c) Paul Denny, James Prather, Brett A Becker, James Finnie-Ansley, Arto Hellas, Juho Leinonen, Andrew Luxton-Reilly, Brent N Reeves, Eddie Antonio Santos, and Sami Sarsa. 2023c. Computing Education in the Era of Generative AI. arXiv preprint arXiv:2306.02608 (2023).
- Donlevy (2023) Katherine Donlevy. 2023. Harvard to roll out AI professors in flagship coding class for fall semester. https://nypost.com/2023/06/30/harvard-to-roll-out-ai-professors-in-flagship-coding-class-for-fall-semester/
- Finnie-Ansley et al. (2022) James Finnie-Ansley, Paul Denny, Brett A Becker, Andrew Luxton-Reilly, and James Prather. 2022. The robots are coming: Exploring the implications of openai codex on introductory programming. In Proceedings of the 24th Australasian Computing Education Conference. 10–19.
- Finnie-Ansley et al. (2023) James Finnie-Ansley, Paul Denny, Andrew Luxton-Reilly, Eddie Antonio Santos, James Prather, and Brett A Becker. 2023. My AI Wants to Know if This Will Be on the Exam: Testing OpenAI’s Codex on CS2 Programming Exercises. In Proceedings of the 25th Australasian Computing Education Conference. 97–104.
- Hellas et al. (2023) Arto Hellas, Juho Leinonen, Sami Sarsa, Charles Koutcheme, Lilja Kujanpää, and Juha Sorva. 2023. Exploring the Responses of Large Language Models to Beginner Programmers’ Help Requests. In Proceedings of the 2023 ACM Conference on International Computing Education Research V.1.
- Kazemitabaar et al. (2023) Majeed Kazemitabaar, Justin Chow, Carl Ka To Ma, Barbara J Ericson, David Weintrop, and Tovi Grossman. 2023. Studying the effect of AI Code Generators on Supporting Novice Learners in Introductory Programming. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–23.
- Kazemitabaar et al. (2024) Majeed Kazemitabaar, Runlong Ye, Xiaoning Wang, Austin Zachary Henley, Paul Denny, Michelle Craig, and Tovi Grossman. 2024. CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24).
- Leinonen et al. (2023a) Juho Leinonen, Paul Denny, Stephen MacNeil, Sami Sarsa, Seth Bernstein, Joanne Kim, Andrew Tran, and Arto Hellas. 2023a. Comparing Code Explanations Created by Students and Large Language Models. In Proceedings of the 2023 Conference on Innovation and Technology in Computer Science Education V. 1. ACM, 124–130.
- Leinonen et al. (2023b) Juho Leinonen, Arto Hellas, Sami Sarsa, Brent Reeves, Paul Denny, James Prather, and Brett A Becker. 2023b. Using Large Language Models to Enhance Programming Error Messages. In Proceedings of the 54th ACM Technical Symposium on Computer Science Education V. 1. ACM.
- Li et al. (2022) Yujia Li, David Choi, Junyoung Chung, Nate Kushman, Julian Schrittwieser, Rémi Leblond, Tom Eccles, James Keeling, Felix Gimeno, Agustin Dal Lago, et al. 2022. Competition-level code generation with alphacode. Science 378, 6624 (2022), 1092–1097.
- Liffiton et al. (2023) Mark Liffiton, Brad Sheese, Jaromir Savelka, and Paul Denny. 2023. CodeHelp: Using Large Language Models with Guardrails for Scalable Support in Programming Classes. arXiv:2308.06921 [cs.CY]
- Liu et al. (2024) Rongxin Liu, Carter Zenke, Charlie Liu, Andrew Holmes, Patrick Thornton, and David J. Malan. 2024. Teaching CS50 with AI: Leveraging Generative Artificial Intelligence in Computer Science Education. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 2 (Portland, OR, USA) (SIGCSE 2024). Association for Computing Machinery, New York, NY, USA, 1927.
- MacNeil et al. (2023a) Stephen MacNeil, Joanne Kim, Juho Leinonen, Paul Denny, Seth Bernstein, Brett A Becker, Michel Wermelinger, Arto Hellas, Andrew Tran, Sami Sarsa, et al. 2023a. The Implications of Large Language Models for CS Teachers and Students. In Proc. of the 54th ACM Technical Symposium on Computer Science Education, Vol. 2.
- MacNeil et al. (2023b) Stephen MacNeil, Andrew Tran, Arto Hellas, Joanne Kim, Sami Sarsa, Paul Denny, Seth Bernstein, and Juho Leinonen. 2023b. Experiences from using code explanations generated by large language models in a web software development e-book. In Proceedings of the 54th ACM Technical Symposium on Computer Science Education V. 1. 931–937.
- Markel and Guo (2021) Julia M Markel and Philip J Guo. 2021. Inside the mind of a CS undergraduate TA: A firsthand account of undergraduate peer tutoring in computer labs. In Proceedings of the 52nd ACM Technical Symposium on Computer Science Education. 502–508.
- Microsoft (2024) Microsoft. 2024. Azure OpenAI. https://azure.microsoft.com/en-us/products/ai-services/openai-service/
- Mirhosseini et al. (2023) Samim Mirhosseini, Austin Z Henley, and Chris Parnin. 2023. What is your biggest pain point? an investigation of cs instructor obstacles, workarounds, and desires. In Proceedings of the 54th ACM Technical Symposium on Computer Science Education V. 1. 291–297.
- Mitra et al. (2024) Chancharik Mitra, Mihran Miroyan, Rishi Jain, Vedant Kumud, Gireeja Ranade, and Narges Norouzi. 2024. Elevating Learning Experiences: Leveraging Large Language Models as Student-Facing Assistants in Discussion Forums. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 2 (Portland, OR, USA) (SIGCSE 2024).
- Ofgang (2023) Erik Ofgang. 2023. What is Khanmigo? The GPT-4 learning tool explained by Sal Khan. Tech & Learn (2023).
- Pardos and Bhandari (2023) Zachary A Pardos and Shreya Bhandari. 2023. Learning gain differences between ChatGPT and human tutor generated algebra hints. arXiv preprint arXiv:2302.06871 (2023).
- Piech et al. (2012) Chris Piech, Mehran Sahami, Daphne Koller, Steve Cooper, and Paulo Blikstein. 2012. Modeling how students learn to program. In Proceedings of the 43rd ACM technical symposium on Computer Science Education. 153–160.
- Prather et al. (2023) James Prather, Brent N Reeves, Paul Denny, Brett A Becker, Juho Leinonen, Andrew Luxton-Reilly, Garrett Powell, James Finnie-Ansley, and Eddie Antonio Santos. 2023. ” It’s Weird That it Knows What I Want”: Usability and Interactions with Copilot for Novice Programmers. arXiv preprint arXiv:2304.02491 (2023).
- Reeves et al. (2023) Brent Reeves, Sami Sarsa, James Prather, Paul Denny, Brett A Becker, Arto Hellas, Bailey Kimmel, Garrett Powell, and Juho Leinonen. 2023. Evaluating the Performance of Code Generation Models for Solving Parsons Problems With Small Prompt Variations. In Proceedings of the 2023 Conference on Innovation and Technology in Computer Science Education V. 1. 299–305.
- Sarsa et al. (2022) Sami Sarsa, Paul Denny, Arto Hellas, and Juho Leinonen. 2022. Automatic generation of programming exercises and code explanations using large language models. In Proceedings of the 2022 ACM Conference on International Computing Education Research-Volume 1. 27–43.
- Sheese et al. (2024) Brad Sheese, Mark Liffiton, Jaromir Savelka, and Paul Denny. 2024. Patterns of Student Help-Seeking When Using a Large Language Model-Powered Programming Assistant. In Proceedings of the 26th Australasian Computing Education Conference. 49–57.
- Soloway et al. (1994) Elliot Soloway, Mark Guzdial, and Kenneth E. Hay. 1994. Learner-centered design: the challenge for HCI in the 21st century. Interactions 1 (1994), 36–48.
- Suzuki et al. (2017) Ryo Suzuki, Gustavo Soares, Andrew Head, Elena Glassman, Ruan Reis, Melina Mongiovi, Loris D’Antoni, and Bjoern Hartmann. 2017. Tracediff: Debugging unexpected code behavior using trace divergences. In 2017 IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC). IEEE, 107–115.
- Wang et al. (2024) Sierra Wang, John Mitchell, and Chris Piech. 2024. A Large Scale RCT on Effective Error Messages in CS1. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 1 (Portland, OR, USA) (SIGCSE 2024). 1395–1401.
- Wu et al. (2022) Yu-Chieh Wu, Andrew Petersen, and Lisa Zhang. 2022. Student Reactions to Bots on Course Q&A Platform. In Proceedings of the 27th ACM Conference on on Innovation and Technology in Computer Science Education Vol. 2. 621–621.
- Yuan et al. (2023) Zhiqiang Yuan, Junwei Liu, Qiancheng Zi, Mingwei Liu, Xin Peng, and Yiling Lou. 2023. Evaluating instruction-tuned large language models on code comprehension and generation. arXiv preprint arXiv:2308.01240 (2023).
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Heather Wei, Amy Xiao, Kitty Gu, Grace Jung, Matthew G Lee, Bjoern Hartmann, and Qian Yang. 2023. Herding AI Cats: Lessons from Designing a Chatbot by Prompting GPT-3. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (Pittsburgh, PA, USA).